OpenStudy (anonymous):

Is this trinomial a trinomial square? m^2+4m-4

OpenStudy (mosaic):

Complete the square and see if it is a perfect square.

OpenStudy (anonymous):

And exactly how do I do that?

OpenStudy (mosaic):

m^2+4m-4 = (m+2)^2 - 2^2 - 4 = (m+2)^2 - 8. Therefore, it is not a perfect square.